### v1

# UINav: A maker of UI automation agents

Abstract
An automation system that can execute natural language instructions by driving the user interface (UI) of an application can benefit users, especially when situationally or permanently impaired. Traditional automation systems (manual scripting, programming by demonstration tools, etc.) do not produce generalizable models that can tolerate changes in the UI or task workflow. Machine-learned automation agents generalize better, but either work only in simple, hand-crafted applications or rely on large pre-trained models, which may be too computationally expensive to run on mobile devices. In this paper, we propose UINav, a demonstration-based agent maker system. UINav agents are lightweight enough to run on mobile devices, yet they achieve high success rates with a modest number of task demonstrations. To minimize the number of task demonstrations, UINav includes a referee model that allows users to receive immediate feedback on tasks where the agent is failing to best guide efforts to collect additional demonstrations. Further, UINav adopts macro actions to reduce an agent’s state space, and augments human demonstrations to increase the diversity of training data. Our evaluation demonstrates that with an average of 10 demonstrations per task UINav can achieve an accuracy of 70% or higher, and that with enough demonstrations it can achieve near-perfect success rates on 40+ different tasks.
**footnotetext: Worked as an intern at Google Research
1 Introduction
The next frontier in artificial intelligence are agents that can autonomously operate computers as humans do. Instructed by users in natural language, these agents are especially valuable when their users have visual or motor disabilities, when they are situationally impaired (e.g., driving, cooking), or when they are engaged in learning new software features and can benefit from in-context guidance. Towards this goal, recently, we have seen a growing interest in agents that fulfill human tasks by manipulating UI elements on a screen Adept (2023), by invoking service APIs Nakano et al. (2021); Komeili et al. (2022), or by interacting with software tools Parisi et al. (2022); Schick et al. (2023). Our work falls in the first category, UI automation, where we propose a demonstration-based agent building system designed to produce lightweight agents that can execute tasks specified in natural language with near perfect success rates.
Existing approaches to UI automation range from manual scripting to large Transformer-based models. Manually-coded UI scripts can work reliably, but they are expensive to write and maintain. This is what enterprises currently use to automate their workflows UIPath (2023); Microsoft Power Automate (2023); Automation Anywhere (2023). Programming by demonstration systems provide a more natural interface, where a user can record a sequence of UI interactions and replay them when needed Kundra (2020); Barman et al. (2016); Riva and Kace (2021); Li et al. (2017). However, these systems do not produce generalizable models that can tolerate changes in user interfaces and workflows, thus leading to high maintenance costs. On the other hand, machine learning approaches to UI automation aim to scale to many tasks and applications. Some previous approaches use imitation learning and reinforcement learning Liu et al. (2018); Gur et al. (2018). However, their testing environments Shi et al. (2017) are much simpler than real-world websites. Moreover, their reward model either needs to be manually programmed or requires large numbers of demonstrations Li and Riva (2021); Humphreys et al. (2022). More recent work Li et al. (2020a); Li and Li (2023); Venkatesh et al. (2022); Wang et al. (2023) leverages the power of Transformers and pre-trained large language models (LLMs) such as PaLM Chowdhery et al. (2022) and GPT-3 Brown et al. (2020). Despite the improvement in performance, these models incur high resource costs (multiple days of training on hundreds of GPUs/TPUs and inference costs beyond the capability of modern devices) which may not always be affordable.
Trade-offs are required between accuracy, generalizability and computational costs when selecting an approach to UI automation. We aim to find a sweet spot between these three properties. We propose UINav, an agent maker system designed to produce lightweight neural agents that can run on mobile devices while yielding high success rates. With UINav, users record task demonstrations and quickly train neural agents to test in the wild.
As in prior work, UINav needs to address the fundamental challenge of how to achieve acceptable success rates with a small number of demonstrations. We observe that the number of demonstrations required to achieve good performance differs widely across tasks and execution environments. If the environment is relatively static even a handful of demonstrations is sufficient to learn. On the other hand, for tasks that have to work across many different UIs more demonstrations may be needed. UINav allows users to quickly and continuously evaluate the performance of intermediate versions of their agents with new tasks and in new environments until a desired quality is achieved. They can do so by using the UINav referee model. This model is trained with the same set of demonstrations used to train the automation agent, but with a different goal, i.e., predicting whether a task is successfully completed (rather than predicting which UI elements to interact with). When testing their agents, users receive immediate feedback on which tasks are failing and may benefit from additional demonstrations.
Another challenge UINav addresses is how to increase the robustness of automation agents against system delays and changes in the UI. It does so through two key techniques. First, every UI action is executed as a small program composed of lower level operations with status checks. These programs, referred to as macro actions, abstract the system-specific details thus greatly reducing the agent’s state space and therefore the number of required demonstrations. Second, UINav adopts demonstration augmentation where human demonstrations are augmented by randomizing non-critical UI elements to increase their diversity.
We develop UINav using an internal dataset of 40+ tasks. We also evaluate it on a public dataset, where UINav outperforms state-of-the-art systems by large margins. Overall, this work makes the following contributions:
  • A practical system designed to build UI automation agents from scratch that achieve near perfect success rates on previously seen tasks and that can be deployed to mobile devices.
  • An error-driven process to collect demonstrations paired with augmentation techniques and macro actions which overall minimize the number of required human demonstrations and increase robustness.
  • A comprehensive evaluation demonstrating UINav’s advantages over state-of-the-art systems.
2 Related work
2.1 UI automation scripts
UI automation scripts consist of low-level action statements (click, type, enter, etc.) each referring to a specific UI element (button, input field, image, etc.). Record-and-replay tools like Selenium Kundra (2020) can be used to facilitate the generation of automation scripts. These scripts can also be wired together and interleaved with API calls to obtain full automation workflows, using robotic process automation tools UIPath (2023); Automation Anywhere (2023); Blue Prism (2023).
To make automation accessible also to non-technical users, programming by demonstration tools have been proposed Sugiura and Koseki (1998); Leshed et al. (2008); Lin et al. (2009); Li et al. (2010); Barman et al. (2016); Li et al. (2017); Chasins et al. (2018). These are advanced record-and-replay tools that can generate fully functional automation scripts and even action graphs Riva and Kace (2021) from recordings of task interactions (demonstrations). Some work also explores whether these demonstrations can be provided in the format of video recordings Bernal-Cárdenas et al. (2020); Chen et al. (2022). Overall, a major downside of this line of work is that these systems do not produce models that generalize to new task workflows and UIs. Every new task and every variation to an existing task requires new demonstrations or manual updates.
To limit the collection of new demonstrations but still provide useful automation, other work examines this problem from an interaction standpoint. PUMICE Li et al. (2019) enables users to resolve ambiguities in their natural language instructions by collaborating with the agent by introducing new concepts and learning which actions the agent supports. SOVITE Li et al. (2020b) displays the system’s understanding of user intents using UI screenshots, thus allowing users to identify failures of the system in understanding their instructions and to correct them using direct manipulation on the screenshots. Unfortunately, these approaches incur an overhead on the user side and make the interaction with the agent less fluid. Moreover, their ability of learning from prior user feedback is limited or completely absent.
2.2 ML-based automation
ML-based automation systems promise to reduce the overhead of UI automation scripts by abstracting single traces into general models. Transformer-based architectures Li et al. (2020a); Bai et al. (2021); He et al. (2021); Banerjee et al. (2022); Li and Li (2023) and reinforcement learning approaches Liu et al. (2018); Gur et al. (2018); Li and Riva (2021) have been proposed to train agents capable of navigating apps and websites when provided with natural language instructions. Yet, it is unclear how well these systems perform in a variety of real-world environments and scale across task categories. Reinforcement learning agents have been tested in synthetic webpages containing only 10–50 UI elements Shi et al. (2017). Existing datasets are mostly for Android: PixelHelp Li et al. (2020a) covers 4 apps for a total of 187 instructions; MoTIF Burns et al. (2022), which we use in this paper, consists of 6,100 episodes across 125 Android apps. As these datasets are collected using app crawling infrastructures and costly human labeling, they make large-scale training difficult.
Recent work has sought to leverage large language models such as PaLM and GPT3 to map natural language instructions into step-by-step instructions (parsing phase) which can be executed in the UI by a simpler grounding model (grounding phase). An example of such a system is UGIF Venkatesh et al. (2022). Wang et al. Wang et al. (2023) also prompt LLMs to perform conversations with a mobile UIs. A main downside of these approaches is the large training overhead (e.g., multiple days of training on hundreds of GPUs/TPUs) as well as the cost of running inference, which is in general too expensive to run on mobile devices.
In this paper, we extend our previous work Li (2021) where macro actions were introduced but was limited to work with OCR and icon recognition, into a full system, that bridges the gap between programming by demonstrations and ML systems by providing an easy-to-learn system to train robust, multi-task AI agents for UI navigation in real-world applications. While the system requires manual demonstrations for training, it provides an error-driven collection of demonstrations where testing scenarios are automatically generated and evaluated by the system, thus reducing the number of redundant demonstrations. The error driven demo collection of UINav is inspired by the DAGGER Ross et al. (2011) algorithm and we show that it is effective in reducing the number of demonstrations for both sequential (UINav referee) and non-sequential (UINav agent) models.
In the next sections we introduce the UINav system. We start by providing some concrete examples of the challenges the system must face with motivates its focus on (i) reducing the number of user demonstrations but also (ii) making demonstrations as diversified as possible.
3 Challenges in UI automation
In this paper we study how UI automation systems can generalize to new execution environments, including different apps and different states in the app, without requiring an excessive number of demonstrations. To illustrate the challenges involved we use an apparently simple task, search, i.e., operating the search bar of an application. The task typically involves typing some search terms into a text box and clicking a search button or pressing Enter.
Two aspects make this task challenging. First, search is a universal task that must work across a large variety of apps where the search bar can take different formats. Some search bars require the user to click on a search icon (typically on the right hand-side) after typing a search term, while others display search results as the user types in; typing in a search bar can automatically open a drop-down menu with autocomplete suggestions or matching products with pictures and names; other search bars have an additional field (e.g., category) typically displayed on the left side of the search bar which must be set beforehand; there are also search bar that are hidden and reveal only after the user clicks on the search icon; etc.
The second axis of complexity regards the start state. When the user initiates a search, the environment may be displaying a page with no search functionality or it may be in an app that is not the target of the search command. In these cases, the first problem the agent has to solve is how to navigate to the state offering the search function, which may involve navigating back in the current app, launching a different app, dismissing welcome screens or ads, etc. Even when the environment already displays the desired search widget, its state may need to be reset, by deleting search terms entered in a previous search or changing the search category. Figure 1 shows the UINav agent searching in YouTube. The agent dismisses popups twice (a) and (b) to reveal the search bar. It then clicks the X button to erase the previous search phrase “something” (c). The system does not reach the desired start state for a search until the screen shown in (d), where the agent sets the focus on the search bar to then enter the search term.
Figure 1: UINav agent searches in YouTube. The pink arrows highlight agent’s actions that are annotated by red boxes and texts. (a) Clicks the back button to dismiss a popup ads; (b) Clicks "X" to dismiss the install page of Audible; (c) Clicks "X" to erase the previously entered search phrase “something”; (d) Focus on the search bar to enter a new search term.
In general, in a real execution environment, an agent is exposed to many different screen conditions caused by a combination of factors, such as different apps, dynamic contents of the apps, previous interactions, layout variance on different devices, UI changes across different app versions and operating systems, ads and notifications, etc. An agent needs to have the capability of ignoring irrelevant UI elements and navigating to relevant states.
One way to model this variability is through more demonstrations, but with a clear overhead. UINav’s first contribution is to adopt an error-driven approach to reduce the number of required demonstrations (§4.2). Its second contribution is to amplify the learning brought by each demonstration by automatically augmenting them (§5.2). Finally, to addresses variability issues due to system delays rather than relying on demonstrations, UINav take a programmatic approach by introducing macro actions (§5.1).
4 System design
Figure 2: Agent and referee development loop and UINav high-level architecture.
Figure 2 illustrates the development workflow and high level architecture of UINav. Given a collection of tasks, each represented by a natural language utterance (e.g., "Add Max Smith to my contacts") and an observation of the current state (e.g., a representation of what is currently displayed on the device’s screen such as the home page or an application page), a neural network-backed agent responds with its choice of action to complete the task. The predicted action is executed by the environment by directly interacting with a device’s system (an emulator or a real phone). Then, the agent is provided with a new observation describing the new state and a new action is predicted. This setup is similar to that of standard reinforcement learning agents, but in UINav the environment also includes a second agent called referee, which is also backed by a neural network. The referee is responsible for judging the completion status of a task (also called episode) at each time step.
In developing UINav agents (left of Figure 2), first, some human demonstrations completing the target tasks are collected. Then, the neural networks of the agent (§4.1.2) and referee (§4.2) are trained (Both networks are initialized with random weights). Next, the trained agent and referee are evaluated on the device system. Failures of either the agent or the referee are recorded and used to guide the collection of new demonstrations which can be used in the next round of training. The development loops over these steps until no more errors of either the agent or the referee are found.
4.1 Agent’s neural network architecture
We now describe the architecture of the agent model (Figure 3). We focus the discussion particularly on the techniques that enable this architecture to reduce the number of required human demonstrations.
4.1.1 Screen representation
An agent receives observations describing the current state of the device, which is represented by the screen currently displayed. A screen consists of a collection of UI elements, with each element described by a set of attributes: type (button, icon, etc.), text description (visible text, content description, resource identifier, etc.), position, utterance matching (whether on-screen text matches words in the utterance, see §4.3), and state (e.g., whether a checkbox is selected). The screen representation can be generated from raw pixels processed by screen understanding techniques Chen et al. (2020); Wu et al. (2021); Zhang et al. (2021), which also include icon detection and text recognition, or from a tree-structured representation of the UI, such as the Android Accessibility View Hierarchy. Our implementation dynamically switches between the two sources of screen representation based on simple heuristics, such as whether the target app is known to provide poor accessibility support or whether the number of accessibility nodes is extremely small (see §7 for more details).
Figure 3: The neural network architecture of the agent model.
4.1.2 Agent model
Figure 3 shows the architecture of the agent model. The input to the neural network is a set of UI elements and an utterance. Each UI element is represented by a vector concatenated from the feature vectors of its attributes. Text labels of UI elements are encoded by a language model such as BERT Devlin et al. (2019). The feature vectors of the UI elements are fed into a Transformer encoder. The output of the Transformer encoder is a function of the encoding of each UI element plus its attention over all other UI elements on the screen, including itself. The utterance is encoded by the language model after its variable sub-strings are masked out (see §4.3 for details).
Given these inputs the agent predicts which action to perform next. This involves predicting: the UI element on which to perform the action, the type of action, and any argument for the action. Our decoder uses a single cross-attention module, with the embedding of the masked utterance serving as the query vector and element encodings from the encoder serving as keys and values. The largest attention weight is used to select the element to act upon, while the vector output of the cross attention module is passed through two independent multi-layer perceptrons (MLP) to predict action type and argument. Together, the encoder and decoder represent a lightweight Transformer model.
In its essence, the agent’s neural network is a scoring system. For any given screen, all the UI elements are scored, and the one with the highest score is selected. Due to the attention in the encoder, for any UI element, its relationship with all the other UI elements on screen can be encoded. Intuitively, the transformer model learns how different combinations of UI elements, and commands map to actions, and uses this knowledge to rank elements to act on. For UI screens it is particularly important that the model learns to evaluate single UI elements in the context of others because the meaning of UI elements is often context sensitive Banerjee et al. (2022) – elements with similar appearance (color, size and shape) can have very different functions but neighboring elements such as text labels or other UI controls can help resolve the ambiguity. For more evidence on how UINav effectively learns to contextually evaluate UI elements, please find some examples in §A.1.2 and §A.1.3.
4.1.3 Action space
The types of action the agent can predict define its action space, summarized in Table 1. Actions can be of two categories. Element actions manipulate a specific element. Global actions are general operations or wrappers for platform-specific functions (e.g., for launching an app). All the tasks that we have tested so far are solvable by these two categories of actions. In the future, we expect to expand the action space to incorporate additional functionality including deep-links and APIs.
Table 1: UINav action space.
Element click <elem>
Clicks the center of the element.
actions focus_and_type <elem,text>
Sets focus on the element, types the text, and optionally presses Enter.
dismiss <elem>
Clicks outside of the element.
Global wait
Waits until the next observation.
actions back
Goes back to the previous screen.
scroll <left|right|up|down>
Scrolls in the specified direction.
open_app <app_name>
Launches the specified application.
4.2 Referee model
In the action space there is no explicit “done” action. This means that a UINav agent does not stop on its own but instead relies on the environment to terminate a task under appropriate conditions. This is common practice in reinforcement learning. Instead of building task-specific termination logic in the environment, we train a referee model to predict whether a task is completed at each step and what its outcome is. The referee model is trained using the exactly same set of demonstrations as the agent, hence it does not incur extra efforts in data collection.
However, the referee also serves a second purpose. A well-known challenge in demonstration-based systems is that they may require excessive developer time to provide a sufficient number of demonstrations Lau (2009) or that it may be difficult to provide samples that are sufficiently different from each other Myers and McDaniel (2001); Lee et al. (2017). By continuously evaluating the execution of a currently-trained agent and identifying failing tasks, the referee guides users towards collecting new demonstrations only for critical scenarios, thus reducing the number of manual demonstrations.
Figure 4: The architecture of the referee model.
The referee is a recurrent neural network (RNN)-based model ( Figure 4). The attention over Transformer-encoded UI elements is similar to that of the agent model, except that the query is the input utterance concatenated with the action history (the action performed in the previous step and its outcome). Although action history could be derived from previous screen representations, feeding it as input directly makes it less challenging as the referee does not have to learn it. The output of the attention module is then fed into a gated recurrent unit (GRU) Cho et al. (2014). The GRU takes this along with the previous internal hidden state as inputs to predict the current status of the step: (1) SUCCESSFUL: the task is completed and it is successful; (2) FAILED: the task has failed or has reached the maximum number of allowed steps; (3) PENDING: the task is ongoing; or (4) INFEASIBLE: the task cannot be executed (e.g., the task may not be well defined). Failed executions are saved along with all their parameters and passed to the demonstrator (see Figure 2).
In the UINav workflow, new human demonstrations are collected only in scenarios where the current version of the agent or the referee make errors. The demonstration collection interface is integrated with the agent and referee. At each step, the agent’s choice of an action and its optional argument are assigned to the internal states and are visualized on the GUI. It is not uncommon that an agent produces correct outputs for unseen scenarios due to the neural networks’ capability of generalization. In such cases, a demonstrator simply proceeds with a single click to the next step, thus avoiding the effort of manually specifying the action parameters. Error-driven demonstration collection significantly reduces human effort as well as the number of training samples, which ultimately leads to lower training times.
4.3 Task workflows
UINav’s focus is on generalizing to different execution environments without requiring an excessive number of demonstrations. However, another large source of variability is the input instruction provided in natural language. To address this problem, as in prior work Liu et al. (2018), we design UINav agents to learn general task workflows rather than specific task instances. We do so by pre-processing utterances to identify sub-string that represent the variables of a task. For example, in Search for tiktok in Google, tiktok is the phrase to search for and can be replaced by other keywords. The variable sub-strings are masked and replaced by placeholders before being encoded, so that the utterance embedding is independent on the specific instances. As a result, there is no need to train with different utterances covering the distribution of variables.
For any utterance, all the replaced sub-strings are included in the list of entities associated with the task. A matching vector is computed for each UI element on the screen and is included in the element attributes passed as input to the agent (see §4.1.1). In the matching vector, each scalar is in the range of $[0,1]$ and computed as the cosine similarity between the text label of the UI element and the corresponding entity string.
Variable sub-strings can be identified by either following pre-defined patterns, through the use of explicit delimiters, or semantic parsers Kamath and Das (2019). LLMs can also be employed Shin and Van Durme (2022); Drozdov et al. (2022); Mekala et al. (2022). UINav still works without utterance masking but may require more demonstrations to reach similar accuracy (see ablation analysis in Table 4).
5 Increasing robustness and efficiency
So far we have illustrated how UINav balances automation accuracy and number of required demonstrations. Now we describe the techniques that increase the robustness of UINav agents in the face of system delays, changes in the UI, and variations in task descriptions.
5.1 Action validation and macro actions
Manipulating UI elements of an actual device involves dealing with various system-specific issues. There are unavoidable delays between the time a state is collected from a device and when an action predicted by the agent is ready to be performed. Screens can also be slow at loading or updating, hence an agent needs to wait for them to stabilize. These delays are particularly noticeable on a mobile device. To deal with these issues, rather than modeling this variability through more demonstrations, UINav takes various programmatic measures .
5.1.1 Action validation
Before executing a predicted action, UINav checks whether a referenced UI element is still on the current screen and if so, whether it has changed. In this way, UINav can detect whether the predicted action is still applicable, and, in case of an obsolete action, requests a new prediction.
5.1.2 Macro actions
In UINav, every action is executed as a small program that is composed of lower level operations with status checks. Such a program is referred to as macro. Macro actions abstract the system-specific details, thus making it possible to build cross-platform agents and simplifying the agent’s logic. Each macro action is implemented following a state transition graph. Figure 5 shows the state transition graph for most macro actions that result in screen changes, such as click and back. It starts at S0, and transitions among the other states according to incoming events, such as Action dispatched and Screen changed, and exits either successfully (S6) or with a failure (S5). The graphs of other macro actions are similar.
Figure 5: The state transition graph for macro actions resulting in screen changes.
Each macro is atomic so that the agent stays idle while a macro is running. During the execution of a macro action, changes to the screen are not visible to the agent, and do not contribute to the state space. In particular, each macro action is designed to encapsulate transitional screens, and finishes when the screen becomes stable or a timeout is reached (required for dynamic screens such as playing a video).
Another advantage of using macro actions is that they package highly dependent, low-level actions. Figure 6 shows an example. The focus_and_type action (inspired from MiniWoB Shi et al. (2017)) consists of 4 low-level actions: clicking the input field to obtain focus, waiting for the blinking cursor to appear, typing the text in the field, and (optionally) pressing Enter. (Note that large arrows in purple are drawn to highlight interesting areas.)
As a result, we are able to utilize a memory-less neural network architecture for the agent. In other words, our agent picks an action based only on the information of the current screen. This makes the neural network easier to train. Additionally, a memory-less neural network can be trained using sets of single screenshots, rather than long sequences of screens which can be hard to collect.
Figure 6: The focus_and_type macro action consists of four steps: (a) clicking the input field (“Search in Drive”) to obtain focus; (b) waiting for the blinking cursor to appear; (c) typing the specified text (“reinforcement learning”); and (d) pressing Enter and wait for the screen to update.
5.2 Demonstration augmentation
To make UINav resilient to UI changes without increasing the number of demonstrations, UINav automatically augments previous demonstrations by randomizing the attributes of randomly-selected, non-critical UI elements. This teaches the agent that these elements may be safely ignored when selecting a suitable action. A UI element is considered critical for a task if it matches one of the following criteria: (1) It is directly manipulated in a demonstration trajectory, (2) Its text label is mentioned in the utterance, (3) Its text label is one of the common phrases appearing on important elements of a UI navigation flow, e.g., $Menu$, $Home$, $Back$, etc., and (4) Its element is of a specific type, such as $Switch$, $Editable$, $Clickable$, etc.
Non-critical UI elements have their attributes modified with a predefined probability by either (i) replacing the embedding of their text labels with random vectors, or (ii) by adding independent random offsets to the four scalars of their bounding boxes, which is equivalent to randomizing both the position and size of UI elements.
6 UINav Console
To collect demonstrations, we have developed a dedicated application, the UINav Console, that can be seen in the right-half side of the screenshots in Figures 1012 in the Appendix. At each step of a demonstration, a user specifies a macro action, including action type, referenced element, and action argument (if any), and then requests execution of the action.
It is typically less effort to complete a task using the UINav Console than directly manipulating the device. For example, entering text using the console takes at most four clicks (clicking the target element, opening the drop-down list of candidate texts, selecting the text to input, clicking the focus_and_type button), while manipulating a real device requires keying-in individual characters. The UINav Console also exposes system APIs, such as opening an app through intents, that are not available through the actual device. While using the console may encourage users to complete a task in a way that is different than how they might do through a native interface, the main goal of a trained agent is to successfully complete tasks. Whether it behaves like a human is less important.
7 Implementation
We built UINav for the Android platform. However, our design is applicable to other platforms and some of our techniques (e.g., macro-actions and screen representation) are specifically designed to be platform agnostic.
Both the agent and the referee models are implemented in TensorFlow. We employ two inference modes, off-device and on-device. During development we use the Python API of TensorFlow to test the models off-device. Once stable, the models are converted to TensorFlow Lite (tflite) for on-device inference. The agent model has about 320k parameters and its tflite version occupies 1.3MB. The referee model has about 430k parameters and it is 1.8MB large. Both models utilize the same pre-trained language model to encode utterances and texts appearing on screens. We choose the smallest model, L-2_H-128_A-2, of SmallBERT Turc et al. (2019), and convert it to a 17.6MB tflite model. Note that no quantization is applied during the tflite conversion of any of the above models. For efficiency, the sentence encoding computation of the agent and referee models are shared.
The neural networks are agnostic to whether the Android Accessibility View Hierarchy or screen understanding techniques are used to produce screen representations. We include demonstrations using both data sources in the same pool of training samples. Both approaches have their limitations. There are icons that are unrecognizable by the icon detectors of screen understanding models and the output of text recognizer may contain errors. On the other hand, visible UI elements may be absent in the corresponding Accessibility tree if the app contains Web views, Canvas, etc.
Table 2 reports the average execution time per task in milliseconds of the agent, referee, and pre-trained BERT models on a high-end and a low-end Android phone. Note that none of the models utilizes any accelerators. Both of the agent and the referee take just a couple of milliseconds on a high-end device and around 5 milliseconds on a low-end device, while the sentence encoding by BERT dominates the total inference time. The time spent in the agent and the referee is stable on different screens, while the sentence encoding by BERT is roughly linear to the number of strings on a screen plus one (for the utterance). We estimate that most of the sentence encoding by BERT can be avoided as there are many texts repeatedly appearing on different screens, and all the variables are masked before encoding any utterance.
Table 2: Average inference times (msec) of the agent, referee, and pre-trained BERT models.
Device Agent Referee BERT Total
High-end phone 1.98 2.21 262.79 267.00
Low-end phone 4.40 5.24 427.63 437.27
For both off-device and on-device modes, we rely on an in-house built companion Android app to extract screen representations and to perform macro actions. For off-device mode, we utilize AndroidEnv Toyama et al. (2021) to communicate between the companion app and our learning environment. For on-device mode, all the models interact with the companion app directly.
8 System evaluation
We evaluate UINav with the following goals in mind: (i) quantifying the number of demonstrations required to achieve high success rates in a variety of tasks, (ii) quantifying the benefits of various techniques we have proposed in comparison to baseline systems, and (iii) evaluating the trade-offs between a multi-task and single-task agent.
Figure 7: Number of demonstrations in the training set grouped by tasks.
Figure 7: Number of demonstrations in the training set grouped by tasks.
8.1 Data collection
To train and evaluate UINav we collect a dataset of 4257 human demonstrations across 43 different tasks and 100+ Android apps and websites. Apps and websites are selected based on popularity (i.e., top apps in the Google Play Store and rank of most visited websites, with the exclusion of apps that do not run on emulators). [^1] Some examples are Amazon, Walmart, Google Messages, Gmail, Contacts, Airbnb, CNN, Zillow, wikipedia.org, linkedin.com, craiglist.org, homedepot.com, target.com, etc. Tasks are decided by enumerating common functionality these apps and websites support. Each task is identified by an utterance pattern. The names of the tasks can be seen along the X axis in Figure 7. Demonstrators collect demonstrations for evaluation independently from those for training. For training, we leverage the UINav error-driven approach to demonstrations, while for evaluation we encourage demonstrators to cover as many task variations as possible. It is not uncommon that there are more evaluation episodes than training episodes for a task. Across all tasks, 3661 episodes are used for training and 596 for testing.
During both demonstration collection and online evaluation, our learning environment performs a few random operations at the beginning of each task, including randomly changing pixel densities, font scales, device orientation, and issuing a sequence of random number of clicks on randomly selected UI elements. The purpose is to start a task from a random start state and to diversify data coverage.
It is unlikely that the agent sees a training sample that is identical to one in the evaluation set due to: 1. the random initialization of a task execution environment; 2. the dynamic characteristics of a live system; and 3) the randomization of feature vectors applied to the UI elements of training samples during demonstration augmentation.
Unless explicitly specified, all the statistics in this paper are for a single pair of agent and referee models trained to complete all the tasks.
8.2 Settings and metrics
We perform two types of evaluation: offline and online. The offline evaluation automatically compares the output of the neural networks with those in an evaluation dataset. It records two metrics: (i) step accuracy: the percentage of task steps where the neural networks and the dataset have matching outputs; and (ii) task accuracy: the percentage of tasks with fully matching steps. The online evaluation interacts with a real phone and checks whether the evaluation task completed successfully. Online evaluation is more precise as there could be more than one way to accomplish a task, and making a mistake in a step may not necessarily result in a failure if the agent recover afterwards. On the other hand, online evaluation of an agent requires human involvement, if the referee is not sufficiently accurate. The results of online evaluation can also vary across sessions due to the dynamic nature of a live system. Hence, we employ the following protocol for online evaluation: each evaluation is composed of 5 sessions, with each session containing 100 randomly sampled task descriptions. We compute the average success rate of each session.
8.3 Accuracy of agent and referee models
Through error-driven demonstration collection, developers can train agents with near perfect success rates (i.e., >99%). Figure 7 shows the number of training samples collected to reach such performance. The SEARCH task required over 1700 task demonstrations because it must work for 100 or more different apps and websites. All other tasks are specific to a single app and thus require fewer samples, 30 on average.
Table 3: Accuracy of agent and referee models (obtained using offline evaluation).
Agent Referee
step acc task acc acc
95.8% 90.6% 99.5%
Table 3 reports the average accuracy of the agent and referee models obtained through offline evaluation, when using all the demonstrations in the training set. The agent’s task accuracy is over 90% and the step accuracy above 95%. The referee model generally exhibits higher accuracy than the agent because its task is simpler (predicting 1 out of 4 labels) and its evaluation is not affected by the ambiguity of multiple potentially correct paths.
8.4 Baseline comparison and ablation analysis
In order to compare with existing UI automation baselines, we train and evaluate UINav agents on the publicly available MoTIF dataset Burns et al. (2022). The dataset includes two test splits: (i) app seen task unseen which tests whether a model can generalize to new tasks, and (ii) app unseen task seen which tests whether a model can generalize to new apps. As in the original MoTIF evaluation Burns et al. (2022), we train UINav using the low-level MoTIF instructions as input utterances, and compare against the same baselines: Seq2Seq Shridhar et al. (2019), MOCA Singh et al. (2020), and Seq2Act Li et al. (2020a).
Table 4: Task and step accuracy on the MoTIF dataset. UINav +aug represents UINav with demonstration augmentation (§5.2) while UINav +aug+utt represents UINav +aug with utterance masking (§4.3).
Model App seen task unseen App unseen task seen
task acc step acc task acc step acc
Seq2Seq Shridhar et al. (2019) 22.5% 40.4% 18.0% 31.3%
MOCA Singh et al. (2020) 21.3% 40.0% 17.0% 32.7%
Seq2Act Li et al. (2020a) 32.4% 66.4% 28.3% 67.7%
UINav 37.9% 73.7% 36.8% 66.8%
UINav +aug 39.4% 74.9% 39.7% 68.4%
UINav +aug+utt 68.3% 89.7% 59.6% 81.9%
Table 4 reports the evaluation results.222‘Task accuracy’ and ‘Step accuracy’ are synonyms of ‘Complete sequence’ and ‘Partial sequence’, respectively, which were used in the MoTIF Burns et al. (2022) and Seq2Act Li et al. (2020a) evaluations. For ablation purposes, we consider three variants of UINav, depending on whether demonstration augmentation (UINav +aug) and utterance masking (UINav +aug+utt) are enabled. The basic UINav model overperforms all three baselines, in all except one case (step accuracy in app unseen and task seen). With demo augmentation (UINav +aug) surpasses all baselines by 7 and 11 percentage points in task accuracy and 8.5 and 0.7 points in step accuracy. This demonstrates the effectiveness of demonstration augmentation in exposing the model to a variety of training conditions which helps it to generalize to new apps and new tasks. Generalizing to new apps is generally harder than generalizing to new tasks, but with the addition of utterance matching UINav achieves 59.6% in task accuracy and 81.9% in step accuracy, well above all baselines.
To evaluate the referee model we use again the MoTIF dataset. The dataset contains task instructions labeled as “feasible” or “infeasible”. Infeasible tasks are those whose step instructions cannot be completed in the given app (e.g., the instruction requires a functionality which is not supported by the app). We compare against the system the authors of the MoTIF dataset propose, which is specifically designed to predict the feasibility or infeasibility of a task. Figure 8 reports the confusion matrix and F1 scores of feasibility/infeasibility predictions. For MoTIF we report results for their best system configuration. [^3] Note that the F1 score reported in their original paper Burns et al. (2022) is incorrect (61.1% instead of 63.5%). The UINav referee predicts 4 states. We map SUCCESSFUL and PENDING to “feasible” and FAILED and INFEASIBLE to “infeasible”. Even though our referee model outputs multiple value states, it produces a significantly better F1 score, 80.9% vs. 63.5%. UINav is especially better in identifying infeasible predictions.
Figure 8: Referee model compared to the MoTIF system Burns et al. (2022) using the MoTIF dataset. For MoTIF we report results for their “best combination”.
8.5 Multi-task vs. single-task agents
By default, UINav agents are multi task, because for on-device applications, it is desirable to keep the total size of all deployed models as small as possible. Our next experiment confirms that this choice is also preferable when the number of training samples is small. We compare the performance of a multi-task agent to that of a single-task agent. We select from our pool the 10 tasks with the largest number of demonstrations. We train one multi-task agent using demonstrations across all 10 tasks and 10 single-task agents using demonstrations from individual tasks. We repeat the training for an increasing number of demonstrations.
As Figure 9 shows, the UINav multi-task agent reaches more than 50% accuracy even with just one demonstration per task. The average accuracy for both multi-task and single-task agents surpasses 80% with 40 demonstrations per task. When the number of demonstrations per task is small, the multi-task agent has a clear advantage, which demonstrates transfer learning across tasks is happening (in line with previous results on MiniWoB Humphreys et al. (2022)). As the number of demonstration per task increases, the accuracy of multi-task agents becomes similar to that of single-task agents, and is slightly higher with over 100 demonstrations per task, probably because the single-task agents suffer less from inter-task interference.
Figure 9: Comparison between multi-task and single-task agents for an increasing number of demonstrations per task.
Figure 9: Comparison between multi-task and single-task agents for an increasing number of demonstrations per task.
9 Limitations
To limit the number of required demonstrations, the UINav agent makes decisions based only on the contents of the current screen and does not utilize information from previous screens. However, if a task truly requires an agent to remember previous states or actions, then the current architecture of the agent model will fail. Our assumption is that a well-designed user interface often presents all the information that is needed for human action on the current screen. Fortunately, the accuracy of our memory-less agent proves that this assumption is true for all the tasks we have tested so far. If in the future, UINav runs into tasks that require memory, the agent model can be enhanced with memory through either a recurrent neural network or padding previous states in its input.
Our approach depends on UI elements for both representing features of screens as well as defining actions. It will not work if a screen representation fails to capture critical UI elements. This can happen also when Accessibility trees miss critical nodes because content embedded in WebViews and Canvas is generally not captured.
Figure 10: The UINav agent sends an email: (a) Clicks the compose button; (b) Types the email address; (c) Types the subject; (d) Types the email content. The action of clicking the send button is not shown due to space limitation.
Figure 11: Two cases of an agent sending a message. The task description is “send the following message in WhatsApp Messenger to Jerry: Are you coming to the meeting?”. (a) In the message view to a different recipient from the one in the utterance; (b) In the message view of the same recipient as the one in the utterance.
Figure 12: An agent selects an action to turn off notification dot (a) when the switch is on, and (b) when the switch is already off. The texts in red (click in a) and (wait in b)) are the actions selected by the agent.
10 Conclusions
In this paper, we presented a demonstration-based system for building UI automation agents capable of performing many tasks in real-world applications. The process we have proposed requires small human effort and no coding skills, while the resulting agents achieve near perfect success rate on previously-seen tasks and also generalize well to new tasks and applications. The trained neural networks are small and fast, thus being suitable for mobile devices.
Acknowledgements
We would like to thank Chris Rawles, Alice Li, Yifan Chang, Marybeth Fair, and Nghi Doan for the help on testing the system.
Appendix A Appendix
A.1 Case study of agent capabilities
In the following figures we report screenshots and the associated UINav console. The large arrows in purple are drawn on the screenshots to highlight interesting areas. In the console it is the annotated screen, where UI elements are identified using blue and green boxes. An element highlighted by a red box indicates that it is selected to receive the next action.
A.1.1 Sending an email with multiple text inputs
Figure 10 shows the image sequence of a UINav agent completing the “send email” task. The task utterance is “send an email to uinav@gmail.com with the subject: Events for the Week and the content: Hi, you are invited to attend our weekly meeting from 4:00pm to 5:00pm on Tuesday (EST - New York)”.
A.1.2 Sending a message to the correct recipient
Figure 11 compares two cases of an agent sending messages. The images are deliberately modified to hide the real names of the recipients. Both (a) and (b) are in the message view of the app but of different recipients, Tom in (a) and Jerry in (b), while the utterance specifies the recipient to be Jerry. The agent correctly recognizes the difference and selects the correct action for both cases: pressing the back button at the top left for (a) and typing the content of the message at the bottom for (b). Note that it is the title bar that contains the information on the current recipient. We believe that it is due to the self-attention of the Transformer encoder that the agent learns whether the text of the title bar matches the recipient is a critical signal in these states.
A.1.3 Understanding the relationship between text label and switch
Figure 12 shows how the UINav agent selects actions to turn off notification dot in two cases: (a) when the switch is on and the agent selects the action to click the text label of "Notification dot on app icon", and (b) when the switch is already off and the agent chooses to wait for the referee to terminate the task. Note that the text label of "Notification dot on app icon" and its switch are independent UI elements in the screen representation, and there are multiple switches on the screen with identical attributes except for their positions and states. The agent learns their relationship probably by the relative positions (horizontally aligned).